OpenStudy (2yoututu11):

The area of a rectangle is 37.5 square inches. The length of the rectangle is 15 inches. What is the width of the rectangle? A. 2 inches B. 2.5 inches C. 22.5 inches D. 562.5 inches

OpenStudy (apoorvk):

L x W = Area or, W = Area/L

OpenStudy (katielong):

it should be 37.5/15 because area = width x height 37.5/15= 2.5 answer is b :)